# Demonstration of sequential processors with quantum advantage and analysis of classical performance limits

### v1

@@ Abstract
In this paper, we theoretically and experimentally analyze sequential processors with limited communication between parts. We compare the expressivity of sequential quantum and classical processors under the same constraints. They consist of three or four modules each of which processes local data. The modules of the quantum processor are linked through one qubit or one qutrit communication, while those of the classical processor communicate through one bit or one trit. For the classical processor, we prove bounds on its performance in terms of inequalities on correlations of the output with a target function. We show that the quantum processor violates these inequalities. We show this violation experimentally on a silicon photonics setup. We describe how to find the classical bound on correlations with arbitrary target function by reducing the problem to minimization of an Ising-type spin-glass Hamiltonian. Our theory is applicable in general problems, such as the low-rank binary matrix approximation.
@@ I Background and overview
Studying and comparing simple structure classical and quantum processors is motivated by the urge to understand the fundamental differences between them. This interest is driven by progress in quantum technology in general and quantum computing in particular. In addition, quick progress in such applications as machine learning raised new questions about the benefits of quantum processing units.
@@
In [1], the authors studied a simple one-qubit processor as an effective classifier. The classical data points $X_{i}$ were encoded in classically tuned parameters of the circuit gates; see Fig. 1. Free parameters of the circuit allowed for training the processor so that the finally measured probability of a chosen state could be used as an indicator of a classification label.
Figure 1: Single qubit and single bit sequential processors of similar structure.
@@
The authors proved the universality of such a classifier, that is, an analogue of the universal approximation theory in a single-layer neural network. This proved that even a single-qubit system can compute complicated functions of parameters of the circuit if it is long enough and is appropriately trained. These ideas were experimentally demonstrated [2, 3, 4, 5].
@@
Universality holds for both a quantum and a classical processors if they are sufficiently extended. In such a situation, a natural question is if there is an advantage of a quantum single-qubit processor over a fairly constructed classical counterpart with fixed finite size. The authors in [1] discussed the possible advantage from entanglement and in [6] from parallelization. In the context of communication complexity [7, 8], the authors formulated a similar problem showing a quantum advantage in sending information in qubit states over the analogous scheme with classical bits. Similar conclusions appeared in the context of temporal entanglement and temporal Bell inequalities [9, 10]. In another paper [11], the author formulated the problem in the context of macro-realism and showed a quantum advantage in speeding-up the computation of certain functions. In [7] an experiment was performed on bulk optics showing that a quantum multiparty communication line with a single qubit of communication (analogous to the sequential processor in Fig. 1) can output a binary target function of local parameters $\{X_{i}\}$ which is not achievable in a single-bit communication classical setup. Also, in [12, 13] the authors showed that certain deterministic outputs from quantum processors cannot be generated by classical circuits in limited space. In this sense, one can speak of a larger expressive power of the quantum processor as in [14, 15].
@@
To show these results, it was necessary to prove rigorous limits on how well some functions could be approximated by the output of classical fixed-structure processors. The functions were shown for which the classical limit was known or established by the authors. However, for arbitrary functions this classical limit is unknown, and a complete understanding of the advantage of the quantum processor over the classical one even in limited space is still a subject of research.
@@
    Is there an advantage of using a fixed length quantum sequential processor with a single-qubit (or single qutrit) communication between processing modules, with respect to the same structure classical processor processing a single bit (or trit)?
@@
    Can we demonstrate it experimentally on the silicon photonic platform?
Our original results are summarized as follows.
  • We report an experimental demonstration of the qubit and qutrit sequential processors on silicon photonics chip and demonstrate the achievable advantage of quantum processors in realizing certain functions of the input parameters. The qubit processor demonstration is analogous to the earlier bulk optics experiment [7], however, different technology–silicon photonics, and different degrees of freedom–a single photon in two modes are involved.
  • To face specific conditions of the design of the silicon photonics chip with four stages and qutrit communication (two photons in two modes) we needed to develop the theory showing the smallest Hamming distance between the deterministic part of the ideal quantum processor output and the optimal output of the corresponding classical processor. We established this bound based on the properties of the quantum processor output and optimization in the space of binary functions.
@@
  • We successfully solved the Ising problems from the previous point for the classical processors with some target functions. Our models involved up to 48 spins with multipartite interaction. The models were solved on a simulated annealing machine Fixstars Amplify.
@@
II Limited communication sequential processor
@@
Consider a sequential processor consisted of several data processing modules, see Fig. 1. In each stage, the module processes the information obtained from the earlier section together with the input provided only locally at this stage $X_{i}$ and unknown to other parts. We assume that the information transferred between the modules is strongly restricted. The last module should output a value that depends on all bits of data $\{X_{i}\}$ inputted at all stages, $O(\{X_{i}\})$. The design of a sequential logic processor consists of optimizing each module so that the output function approaches the desired target function $T(\{X_{i}\})$. The latter can be interpreted as a truth table for a logical function with Boolean variables $\{X_{i}\}$.
@@
In Fig. 1 $a)$ we show a three-module quantum sequential processor with two classical bits $(X_{i},X_{i+1})$ processed in each module and a single-qubit inter-module communication. Fig. 1 $b)$ shows its classical counterpart with a one-bit communication. In the modules of the quantum processor, we allow for unitary transformations as well as measurements and classical functions. Therefore, the quantum processor from Fig. 1 $a)$ is at least as powerful in any task as the corresponding classical processor. In the classical circuit, in addition to the restriction that intermodular communication is limited, there are no additional restrictions on the local functions performed by the modules.
In our research, we consider a specific realization of the quantum processor described in detail in the following sections.
@@ II.1 Quantum one-qubit processor
Consider a specific three-stage quantum processor with single-qubit communication between the modules and two classical bit inputs $(X_{2i},X_{2i+1})$ at each module. Depending on the input bits, we fix one of four one-qubit gates on the qubit obtained from the earlier part:
@@
$(0,1)$ $\rightarrow$ $\hat{\sigma}_{z}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$, (2)
@@
The initial qubit state is $|0\rangle$. In the end, the qubit is measured in the computational basis $\{|0\rangle,|1\rangle\}$. For all $2^{6}$ input bit values $X_{1},...,X_{6}$ ordered in lexicographical order, the final measurement of our idealized processor returns either deterministic outputs $|0\rangle$ or $|1\rangle$, or a state which is a superposition of the two basis states. We define the target function as follows:
  • $T(\{X_{i}\})=1$ if the output is deterministically $|0\rangle$
  • $T(\{X_{i}\})=-1$ if the output is deterministically $|1\rangle$
  • $T(\{X_{i}\})=0$ if the output state is a superposition of $|0\rangle$ and $|1\rangle$.
The target function, after reshaping into a matrix, for convenience, can be written explicitly as follows,
$T(X_{1}X_{2},X_{3}...X_{6})=\left[\begin{array}[]{cccccccccccccccccccc}-1&1&0&0&1&-1&0&0&0&0&1&-1&0&0&-1&1\\ 1&-1&0&0&-1&1&0&0&0&0&-1&1&0&0&1&-1\\ 0&0&1&-1&0&0&-1&1&1&-1&0&0&-1&1&0&0\\ 0&0&-1&1&0&0&1&-1&-1&1&0&0&1&-1&0&0\end{array}\right]$ (5)
where the first two bits determine the row and the remaining four bits determine the column. This notation can be associated with the Karnaugh maps [16] known in the literature on logic design. Alternatively, it can be seen as a tensor as we refer to it in Section VI.
We used this target function to calculate the correlation with the realistic experimental output. The experiment and the results are presented in Section III. There, based on the experimental output $O(x_{i})$ of $N$ randomly selected sequences $x_{i}$ of six bits $X_{1},...,X_{6}$ we calculate the correlation function
@@
In Section IV we prove bounds on the achievable correlation function for optimized classical sequential processor with three stages, and one-bit communication to be compared with the experimental results.
@@ II.2 Quantum one-qutrit processor
We do a similar analysis for a quantum sequential processor with 3 and 4 modules and a communication carried forward by a three-level quantum system (qutrit). Our analysis is tailored to a quantum optics experiment on silicon photonics, so the qutrit we consider is realized by a state of two photons in two modes. The Hilbert space of the system is spanned by three vectors in the photon number notation $\rangle, |02|02\rangle$, and $|11\rangle$, where $|20\rangle$ means that there are two photons in the first mode and zero photons in the second, etc. As input state, we consider $|20\rangle$. As gates, depending on each module input bits $(X_{2i},X_{2i+1})$, we take:
  • For the three-stage processor and for modules 2,3, and 4 of the four-stage processor we applied gates (1)-(4) acting on the photon creation operators of the modes.
  • For module 1 of the four step processor, because the part of the circuit of our chip did not allow to apply the gates as in the other modules, we applied the following gates
@@
If the final measurement of an ideal processor deterministically indicates the state $|20\rangle$ we define the entry of the corresponding target function as 1. If it deterministically indicates the state $|02\rangle$, we define the entry 1-1. Otherwise, the target function entry is set to 0. The quantum processor we built does not allow us to achieve state $|11\rangle$ deterministically from the initial state $|20\rangle$ by any unitary transformation that preserves the bosonic commutation relations. Therefore, the target function of our qutrit processor resembles a qubit processor target.
@@
III Silicon photonics experiment
We implemented both a one-qubit and a one-qutrit processor on a silicon photonic integrated circuit and evaluated their performance using correlation functions [17, 18, 19]. As shown in Fig. 2(a), the experimental setup employed a $\times 4$ universal unitary transformation circuit composed of beam splitters and phase shifters [20, 21]. In this study, we used two spatial modes of the circuit to construct a sequential processor by cascading multiple unitary transformations. Each unitary transformation consisted of a phase shifter and a Mach–Zehnder interferometer (MZI), allowing us to realize the target operations with high precision.
Figure 2: (a) Schematic of the sequential quantum processor implemented on a silicon photonic chip. The unitary gate was implemented using a heater and an MZI enclosed by the red dotted lines. (b) Overview of the one-qubit processor. A heralded single photon was injected into the sequential processor. (c) Overview of the one-qutrit processor. Two photons generated from attenuated laser light at the single-photon level were injected into the sequential processor. In Figs. (b) and (c), the circuit enclosed by the red dotted lines corresponds to a single unitary gate in each case.
For the implementation of the one-qubit processor, heralded single photons were employed, as illustrated in Fig. 2(b). Photon pairs were generated via spontaneous four-wave mixing (SFWM), utilizing the third-order nonlinearity of silicon [22, 23]. Specifically, a continuous-wave (CW) pump light with a wavelength of 1549.32nm and an output power of approximately 10mW was injected into a silicon waveguide, generating signal and idler photons through SFWM. The signal photons had a central wavelength of 1559.0 $\pm$ 0.8nm, while the idler photons had a central wavelength of 1539.8 $\pm$ 0.8nm. The spectral width of these photons is determined by the bandwidth of the off-chip frequency filter (200 GHz). The signal photons were used as heralds (triggers), and the idler photons were directed to the processor circuit [24, 25, 26]. After passing through these filters, both the idler and corresponding signal photons were detected using superconducting nanowire single-photon detectors (SNSPDs). The generation rate of signal single photons obtained after passing through the circuit was approximately 30 counts per second.
In this experiment, the input state was $|1;0\rangle_{ab}$, where a single photon was present in mode $a$. Randomly selected operations from the set $\{\hat{I},\hat{\sigma}_{z},\hat{H},\hat{H}^{*}\}$ were applied through three successive unitary transformations. At the output, the occurrence probabilities of the states $|1;0\rangle_{ab}$ and $|0;1\rangle_{ab}$ were measured from the photon number distribution obtained over a 1-second integration period. Using approximately 85,000 detection events, the correlation function with respect to the target state was calculated.
Fig. 3 shows a histogram of the obtained correlation function. To construct the histogram, the approximately 85,000 experimental data points were divided into subsets of 1,000 events each. The correlation function was calculated for each subset, and the frequency distribution of these values was plotted. From the histogram, the mean of the correlation function was 0.490, and the square root of the variance (standard error of the mean) was 0.017. These results indicate that the experimentally obtained correlation function significantly exceeds the limit of 0.25 achievable by the classical processor with one bit communication. In the figure the vertical blue line corresponds to this limit. Its value is derived in the next section.
Next, for the implementation of the one-qutrit processor, we attenuated a laser beam to the single-photon level and approximated a two-photon state by post-selecting events in which two photons were detected simultaneously (Fig. 2(c)). The total number of two-photon detection counts observed in this experiment was approximately 5,000 per measurement.
The input state for the qutrit experiment was $|2;0\rangle_{ab}$, corresponding to two photons in mode $a$. Four unitary transformations, randomly selected from the operations defined in Eqs.  and (1)–(4) and (7)–(10), were sequentially applied. The output probabilities of the states $|2;0\rangle_{ab}$ and $|0;2\rangle_{ab}$ were measured, and approximately 65,000 detection events were collected. Correlation functions were then calculated based on these data.
Fig. 4 shows a histogram of the correlation function for four module quantum processor with one qutrit communication. As in Fig. 3, the approximately 65,000 experimental data points were divided into subsets of 1,000 events each, and the correlation function was calculated for each subset. The histogram displays the frequency distribution of the resulting correlation values with respect to the target function (11). From the histogram, the mean of the correlation function was 0.525, and the square root of the variance was 0.014. These results indicate that the experimentally obtained correlation function significantly exceeds the classical limit which, as proven in the next section and shown in Fig. 4, is 0.375.
Figure 3: The histogram of the correlation function based on the data from 3-module single qubit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation which for the considered target function is equal to 0.25.
Figure 4: The histogram of the correlation function based on the data from 4-module single qutrit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation which for the considered target function is equal to 0.375.
Figure 5: The histogram of the correlation function based on the data from 3-module single qutrit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation which for the considered target function is equal to 0.4375.
Finally, in post-processing, from the data of the qutrit experiment with four modules we extracted data which correspond to the first gate fixed as the identity. In this way we performed the analysis of the three-module processor with a qutrit communication. The target function in this case is given by the first four rows of (11). We built the histogram of the values or correlation function for 65 subsets of data with approximately 250 points in each subset. The average is 0.509 with the standard error of the average 0.068. This allows us to conclude that the observed value of the correlation significantly exceeds the limit for the correlation achievable by the classical processor with one-trit communication which, as proven in the next section and shown in Fig. 5, is equal to 0.4375.
In conclusion, in each case which we analyzed, we observe that the realistic experiment produces the correlation function close to 0.5, which is the maximum value of the correlation function with respect to the target (5) or (11) with half of the entries equal to zero. In Section IV we establish bounds on the achievable correlation function for optimized classical sequential processors which are shown in Figs. 3, 4, and 5 as the vertical blue lines. We conclude that the classical processor cannot output a function that approaches the target function for any deterministic strategy. In this sense, the quantum processor significantly outperforms the classical one in this task, showing a separation of expressivity between the classical and quantum processors.
IV Classical bounds on correlations
In this section, we establish bounds on the correlations with the target functions given in (5), (11) which are achievable by outputs of sequential classical processors with one-bit and one-trit communication, respectively, as shown in Figs. 3, 5, 4. The maximum correlation is uniquely defined by the minimum number of errors between the output of the classical processor and the target function. In the following parts, we establish these results as propositions.
IV.1 Bound for one-bit processor
Proposition 1.
Proof.
In the proof, we explicitly use the properties of the target function (5). The rows of this matrix correspond to the processor action depending on the first-stage output. As the first module can have only two different symbols, the target reshuffled as a $4\times 16$ matrix needs to be rank 2. So, two lines need to be identical. Let these lines be 1st and 3rd (the same can be repeated for all other choices). We add the two rows, obtaining a binary vector of length 16.
We reshape this vector into a matrix $4\times 4$. It needs to be rank 2, as the rows describe the processor’s actions depending on the output from the second stage, which can output at most two symbols. However, we observe that each row is different and that the minimum number of different bits between each pair of rows (minimum Hamming distance) is 2. We need to correct two bits in two rows. In this way, we necessarily introduce 4 errors.
Finally, we have the remaining lines 2 and 4 of the target reshaped into the matrix $4\times 16$. When we reshape each of these lines into $4\times 4$ matrices, for the same reason as before, we need to correct 4 additional bits.
Performing simulated annealing combinatorial optimization on Fixstars Amplify (details of this procedure are provided in the next section) we find a strategy for each module of the processor such that the final output approximates the non-zero part of the target with exactly 8 errors. The example is shown in Table 1.
$i$ $F_{i-1}=0$ $F_{i-1}=1$
$F_{1}$ (0,1,1,1) -
$F_{2}$ (0,1,1,0) (1,0,0,1)
$F_{3}$ (-1,1,-1,1) (1,-1,1,-1)
Table 1: Optimal strategies for one-bit processor with three nodes.
This concludes the proof. The minimum 8 errors correspond to the maximum value of the correlation between the output and the target equal to 0.25. ∎
Actually, because the size of the problem is relatively small, we could guess the solution with only 8 errors. However, we perform the simulation on the annealer to demonstrate that our numerical optimization tool works well.
IV.2 Bound for three-module trit processor
Proposition 2.
The number of errors between the output of the three-stage processor with one-trit communication and the nonzero part of the target function (11) (lines 1-4) cannot be less than 2.
Proof.
The target function is the first four rows of (11). As the rows of the matrix correspond to the functions in modules 2 and 3 given one of 3 symbols output from module 1, the rank of the matrix has to be equal to 3. Because the matrix is binary, the rank restriction requires that two of the rows be identical.
Without loss of generality, assume that the first and third rows are identical (the procedure can be repeated in all other cases). We complete the undefined elements of the first row by the defined elements of the third row. We observe that we have four blocks of size 4. Moreover, all are different, and the number of different bits between each pair of the four-blocks is equal to 2. We can reduce the number of these 4-size blocks to three by changing at least 2 bits. So, the smallest number of changes we would need is 2.
We can give an explicit example of a processor which approximate the target function with only two errors, as Table 2.
@@ IV.2 Bound for three-module trit processor
IV.3 Bound for four-module trit processor
Proposition 3.
Proof.
The proof uses the properties of the target function (11) explicitly.
We reshape the target into the $4\times 64$ matrix. The rows correspond to the processor action depending on the first-stage output. As it can have only three different symbols, the target reshuffled as a $4\times 64$ matrix must be rank 3. So, two lines need to be identical. Let these lines be the first and second (the same can be repeated for all other choices). We add the two rows, obtaining a binary function of length 64.
We reshape this vector into a matrix $4\times 16$. It needs to be rank 3, as the rows describe the processor’s actions depending on the output from the second stage, which can contain at most 3 symbols. So, two rows of this matrix have to be the same. However, we observe that the minimum number of different bits between each pair or rows (minimum Hamming distance) is 8. Correcting some entries in one of the rows, we necessarily introduce 8 errors in the original target function.
We observe that the remaining lines of the $4\times 16$ matrix contain 4 different blocks of length 4. To reduce this number to 3 different blocks of length 4 we need to introduce 2 additional errors per line. This means that 4 additional errors are required to correct these parts.
Finally, we have the remaining lines 3 and 4 of the target reshaped into the $4\times 64$ matrix. When we reshape each of these lines into $4\times 16$ matrices, for the same reason as before, we expect that each of them is rank 3. So, the two lines need to be identical. We observe that when we add two complementary lines we get vectors of length 16 consisting of 4 blocks of size 4. To reduce this number to 3 we need to introduce 2 errors per each of the two rows.
Performing simulated annealing combinatorial optimization on Fixstars Amplify (details of this procedure are provided in the next section) we find an optimal strategy for each module of the processor such that the final output approximates the non-zero part of the target with exactly 16 errors. The example is shown in Table 3.

### v2

@@ Abstract
In this paper, we theoretically and experimentally analyze sequential processors with limited communication between parts. We compare the expressivity of sequential quantum and classical processors under the same constraints. They consist of three or four modules, each of which processes local data. The modules of the quantum processor are linked through one-qubit or one-qutrit communication, while those of the classical processor communicate through one bit or one trit. For the classical processor, we prove bounds on its performance in terms of inequalities on correlations of the output with a target function. We theoretically show that the quantum processor violates these inequalities. We show this violation experimentally on a silicon photonics setup. We describe how to find the classical bound on correlations with arbitrary target function by reducing the problem to the minimization of an Ising-type spin-glass Hamiltonian. Our theory is applicable in general problems, such as the low-rank binary matrix approximation.
@@ I Introduction
Fast progress in quantum technology in general and quantum computing in particular urges us to identify the features responsible for the fundamental difference between classical and quantum processors. In addition, quick progress in applications such as machine learning has raised questions about the benefits of quantum processing units. Therefore, a rigorous comparison and meaningful benchmarking of classical and quantum processors, even with simple structure, is crucial for further development of the field.
@@
In Ref. [1], the authors studied a simple one-qubit processor as an effective classifier. The classical data points $X_{i}$ were encoded in classically tuned parameters of the circuit gates (see Fig. 1). The free parameters of the circuit allowed for training the processor so that the probability measured in the final for a chosen state could be used to distinguish the correct classification label.
Figure 1: Single-qubit and single-bit sequential processors of similar structure.
@@
The authors proved the universality of such a classifier, that is, an analog of the universal approximation theory in a single-layer neural network. This showed that even a single-qubit system can compute complicated functions of circuit parameters, assuming that this system is long enough and is appropriately trained. These ideas were experimentally demonstrated [2, 3, 4, 5].
@@
Universality holds for both a quantum and a classical processor if they are sufficiently extended. In such a situation, it is natural to ask if there is an advantage of a quantum single-qubit processor over a fairly constructed classical counterpart with fixed finite size. The authors in Ref. [1] discussed the possible advantage from entanglement and in Ref. [6] from parallelization. In the context of communication complexity [7, 8], the authors formulated a similar problem showing a quantum advantage in sending information in qubit states over the analogous scheme with classical bits. Similar conclusions appeared in the context of temporal entanglement and temporal Bell inequalities [9, 10]. In another paper [11], the author formulated the problem in the context of macro-realism and showed a quantum advantage in speeding up the computation of certain functions. In Ref. [7], an experiment was performed on bulk optics showing that a quantum multiparty communication line with a single qubit of communication can output a binary target function of local parameters $\{X_{i}\}$. The same target function is not achievable in a single-bit communication classical setup. The setup of this experiment was similar to that in Fig. 1. In addition, in Refs. [12, 13] the authors showed that certain deterministic outputs from quantum processors cannot be generated by classical circuits in limited space. In this sense, one can speak of a larger expressive power of the quantum processor as in Refs. [14, 15].
@@
To show these results, it was necessary to prove rigorous limits on how well some functions could be approximated by the output of classical fixed-structure processors. For the functions shown, the classical limit was known or established by the authors. However, for arbitrary functions this classical limit is generally unknown. Thus, a complete understanding of the advantage of the quantum processor over the classical processor even in limited space is still a subject of research.
@@
    Is there an advantage in using a fixed-length quantum sequential processor with a single-qubit (or single-qutrit) communication between processing modules, with respect to the classical processor of the same structure processing a single bit (or trit)?
@@
    Can we demonstrate the advantage experimentally with sufficient accuracy on nowadays available platforms? In particular, in this work we focus on a silicon photonic platform.
Our original results are summarized as follows:
  1. 1.
    We report an experimental realization of the qubit and qutrit sequential processors on the silicon photonic chip. We demonstrate quantum processors that achieve an advantage in computing certain functions of the input variables. The demonstration of a qubit processor is analogous to the previous bulk optics experiment [7]. However, it involves different technology, the use of silicon photonics, and different encoding—a single photon in two modes.
  2. 2.
    To face specific conditions of the design of the silicon photonics chip with four stages and qutrit communication (two photons in two modes), we needed to develop the following method. We sought for the smallest Hamming distance between the deterministic part of the ideal quantum processor output and the optimal output of the corresponding classical processor. We established this bound based on the properties of the quantum processor output and optimization in the space of binary functions.
  3. 3.
@@
  4. 4.
    We successfully solved the Ising problems from the previous point for the classical processors with given target functions. Our models involved up to 48 spins with multipartite interactions. The models were solved on a Fixstars Amplify simulated annealing machine.
  5. 5.
@@
II Sequential processor with limited communication
@@
Consider a sequential processor consisting of several data processing units (see Fig. 1). In each stage, the unit processes the information obtained from the earlier section together with the input provided only locally at this stage, $X_{i}$, and unknown to other parts. We assume that the information transferred between the modules is strongly restricted. The last module should output a value that depends on all bits of data $\{X_{i}\}$ inputted at each stage, $O(\{X_{i}\})$. The design of a sequential logic processor consists of optimizing each unit so that the output function approaches the desired target function $T(\{X_{i}\})$. The latter can be interpreted as a truth table for a logical function with Boolean variables $\{X_{i}\}$.
@@
In Fig. 1(a)(a) we show a three-module quantum sequential processor with two classical bits $(X_{i},X_{i+1})$ processed in each stage and a single-qubit inter module communication. Figure 1(b)(b) shows its classical counterpart with, this time, a one-bit communication. In the modules of the quantum processor, we allow for unitary transformations, measurements, and classical functions. Therefore, the quantum processor from Fig. 1(a)(a) is at least as powerful in any task as the corresponding classical processor. In the classical circuit, in addition to the restriction that communication between units is limited, the local functions performed by the modules do not have any additional restrictions.
The goal of our research is to show an advantage of quantum processor with respect to the classical one with similar constraints. Consider the following toy example as an illustration. Assume that we want to design a simple distributed circuit that calculates the sum of several two-bit numbers modulo 4. We also assume that the total sum is promised to be either 0 or 2. We consider that the input numbers are loaded to the circuit in different modules and each of them can only communicate by sending a single bit, or a single qubit, to the next unit. The quantum processor with a single-qubit communication can find the exact solution to the above problem. The scheme is shown in Fig. 2. The strategy is to choose a subset of four points on a large circle on the Bloch sphere. We encode the two-bit numbers $x$ as a $x\pi/2$ rotation acting on the state transmitted by the previous module without measurement. In this way, the quantum processor can successively rotate the state vector between the four points performing addition of two-bit numbers modulo four. The promise that the total sum is in a given one-bit subset (0 or 2) implies the deterministic output of the correct result if the final state is measured in the corresponding basis. This procedure can be applied for any set of numbers satisfying the promise. The correlation computed between the target and the output of the quantum processor is maximal.
For the classical processor with one-bit communication, at each state, one must decide the state of the single bit to be passed to the next module. As a two-bits number is needed to specify the partial sum modulo 4, the classical processor that passes only one bit of information results in a possible error introduced at each stage. Therefore, this processor cannot output the correct solution of the problem for different sets of input numbers with certainty. This causes reduction of the correlation between the output and the target when the input is chosen randomly.
Figure 2: Example. Adding four two-bit numbers $(0,2,3,2)$ modulo 4 with a single-qubit quantum processor. The processor uses operations that rotate a qubit state vector, shown here as a thick violet arrow, between four points, indicated by thin yellow arrows, on the Bloch sphere. The added numbers can be arbitrary two-bit numbers with only one constraint that the total sum modulo 4 is either 0 or 2. For any set of two-bit numbers with the above promise, the ideal quantum processor solves the addition problem without errors.
It is easy to imagine other functions that could be solved by simple quantum processors. For example, adding three-bit numbers modulo 8 with appropriate promise. Although the class of target functions and problems that can be solved by simple or advanced quantum processors is not fully known, we can say that this class is non trivial and contains interpretable examples interesting from a perspective of logic, arithmetic, communication complexity, and fundamental aspects of quantum mechanics [7, 8, 12, 13].
In this paper, we consider deterministic operations with binary targets. Allowing targets to be achievable probabilistically opens yet another space for comparison between classical and quantum processors with fixed constraints. In that case, separation between performances of classical and quantum processors would result in different probability distributions as functions of the input variables. The distributions that would be unachievable by classical processors would lead to violation of variations of Bell-type inequalities established for given constraints. Some classes of such inequalities are well studied. So, the phenomenon of quantum advantage in achieving certain correlations in restricted systems is typical in the sense of not being limited to small number of separated examples. However, studying typicality, which would require different approaches, is not a subject of the present research. We focus on a few target functions tailored to our experimental setup for simplicity. This is enough to prove the existence of an advantage and technical feasibility of its demonstration. It should not be considered as a fundamental restriction. Other setups can be designed to demonstrate similar advantage and general procedures we worked out can be helpful for establishing rigorous results in other cases.
@@ II.1 Quantum one-qubit processor
We consider a specific three-stage quantum processor with single-qubit communication between modules and two classical bit inputs $(X_{2i},X_{2i+1})$ at each module. Depending on the two input bits, we choose one of the four one-qubit gates to be applied to the qubit obtained from the earlier part:
@@
$(0,1)$ $\rightarrow$ $\hat{\sigma}_{x}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$, (2)
@@
The initial qubit state is $|0\rangle$. In the end, the qubit is measured in the computational basis $\{|0\rangle,|1\rangle\}$. For all $2^{6}$ input bit values $X_{1},...,X_{6}$ ordered in lexicographical order, the final measurement of our ideal processor returns either deterministic outputs $|0\rangle$ or $|1\rangle$, or a state which is a superposition of the two basis states. We define the target function as follows:
$T(\{X_{i}\})$ $=1\text{ if the output is deterministically }|0\rangle,$ (5)
$T(\{X_{i}\})$ $=-1\text{ if the output is deterministically }|1\rangle,$ (6)
[…]
The target function, after being reshaped into a matrix, for convenience, can be written explicitly as follows:
$T(X_{1}X_{2},X_{3}...X_{6})=\left[\begin{array}[]{cccccccccccccccccccc}-1&1&0&0&1&-1&0&0&0&0&1&-1&0&0&-1&1\\ 1&-1&0&0&-1&1&0&0&0&0&-1&1&0&0&1&-1\\ 0&0&1&-1&0&0&-1&1&1&-1&0&0&-1&1&0&0\\ 0&0&-1&1&0&0&1&-1&-1&1&0&0&1&-1&0&0\end{array}\right],$ (8)
where the first two bits index the row and the remaining four bits index the column. This notation is similar to the Karnaugh maps [16] known in the literature on logic design. Alternatively, it can be seen as a tensor, and this is how we refer to it in Sec. V.
We used this target function to calculate the correlation with the experimental output with real experimental imperfections. The experiment and results are presented in Sec. III. There, based on the experimental output $O(x_{i})$ of each of $N$ randomly selected sequences $x_{i}$ of six bits $X_{1},...,X_{6}$, we calculate the correlation function
@@
In Appendix A, we prove that the bound on the achievable correlation function for an optimized classical sequential processor with three stages, and one-bit communication is 0.25. This corresponds to the minimum number of errors in the outputs of the optimal classical processor with respect to the non zero part of the target equal to 8. An optimal strategy for this processor is given in Table 1.
$i$ $F_{i-1}=0$ $F_{i-1}=1$
$F_{1}$ (0,1,1,1) -
$F_{2}$ (0,1,1,0) (1,0,0,1)
$F_{3}$ (-1,1,-1,1) (1,-1,1,-1)
Table 1: Optimal strategies for one-bit processor with three nodes. The symbols correspond to the notation from Fig. 1. For example, here $(0,1,1,0)$ are bits communicated by the second node when the local inputs are $(00,01,10,11)$ respectively, and the bit 0 is obtained from the previous node.
The processor described above realizes the adder of two-bit numbers modulo 4 from the previous example. Here, bits $00$ are associated with number 0, and $10$, $01$, and $11$ with 1, 2, and 3, respectively. The zeros of the target function correspond to the situation when the promise from the problem formulation is not satisfied. Target functions with many zeros, as in Eq. (8) are typically considered in logic circuit design, where the truth table is specified only on desired entries. Leaving the remaining entries not specified allows for more freedom in designing the optimal circuit. Therefore, not fully specified target functions should not be considered as a restriction.
@@ II.2 Quantum one-qutrit processor
We perform a similar analysis for a quantum sequential processor with three and four modules and a communication carried forward by a three-level quantum system (qutrit). Our analysis is tailored for a quantum optics experiment on silicon photonics, so the logical state of the qutrit we consider is physically encoded by the state of two photons in two different modes. Each logical state is uniquely described by a physical state in the two-mode Fock basis as follows:
[…]
where $|20\rangle$ states for two photons in the first mode and zero photons in the second, etc.
As the input state, we take $|20\rangle$. As gates, depending on each module input bits $(X_{2i},X_{2i+1})$, we choose the following:
  • For the three-stage processor and for modules 2, 3, and 4 of the four-stage processor, we applied gates [Eqs. (1)-(4)] acting on the photon creation operators in each mode.
  • For module 1 of the four-step processor, the design of the circuit of our chip did not allow for applying the same gates as in the other modules. Instead, we applied the following gates
@@
If the final measurement of an ideal processor deterministically indicates the state $|20\rangle$, we define the entry of the corresponding target function as 1. If it deterministically indicates the state $|02\rangle$, we define the entry 1-1. Otherwise, the target function entry is set to 0. The quantum processor we built does not allow us to achieve state $|11\rangle$ deterministically from the initial state $|20\rangle$ by any unitary transformation that preserves the bosonic commutation relations. Therefore, the target function of our qutrit processor resembles a qubit processor target.
[…]